498 mL of a 0.21 M sodium hydroxide solution is added to 158 mL of a 0.61 M rubidium hydroxide solution. Calculate the pOH of the resulting solution.
Answer #1

498 mL of 0.21M Naoh solution molarity (M)= 0.2189 Volume (u) = 498 mL = 498x10*3 L So, moles of Naon = molarity (M) x Volometotal volume of Solution aftor = (498+ 158 m2 mixing = 656 me so, concentration of Cora) - mous Volume = 200.96 mmol - 656 m2